OpenStudy (anonymous):

Given that F(x) = the integral of the sqrt of (t^2+9) from to 0, calculate F''.

OpenStudy (anonymous):

\[F(x) = \int\limits_{0}^{x} \sqrt{t^2+9}\]

OpenStudy (anonymous):

there shud be a dt after the integral

OpenStudy (anonymous):

F'(x)= \[ \sqrt{t ^{2}=9}\] then find the derivative of that to find F''(x). you'll need the chain rule

OpenStudy (anonymous):

let t = f(x) = x\\[\[\int\limits\limits_{0}^{x} (\sqrt{x^2 + 9} )dx\]\] by the Fundamental Theorem of Calculus, integrals and derivatives just cancel out, so: dx/dy \[dx/dy \int\limits\limits_{0}^{x} (\sqrt{x^2 + 9} )dx = \sqrt{x^2 + 9}\] now just derive that to get F''(x)

OpenStudy (anonymous):

im sorry! substitute that t for x in the F'(x)

OpenStudy (anonymous):

oh right what tmmss5533 said, you'll have to chain rule the x^2 because it really is f(x)^2

OpenStudy (anonymous):

wait so wats the answer???

OpenStudy (anonymous):

wow im terrible lol i meant to put x^(2)+9 under that radical

OpenStudy (anonymous):

is the answer x/ sqrt of (x^2+9)

OpenStudy (anonymous):

Well what's the derivative of x^2? It just 2(x) right? so 2(x) d/dx(sqrt(x^2 + 9) whats the derivative of a sqrt(x^2 + 9)? It's gonna be 1/sqrt(x^2 + 9) so putting them together: 2x/[sqrt(x^2 + 9)]

OpenStudy (anonymous):

can u help me with anotehr problem

OpenStudy (anonymous):

Hermeezey can you please help me with my last two recent math questions?

OpenStudy (anonymous):

lol i asked hermezy first

OpenStudy (anonymous):

LOL post your problems and i'll see what i can do

OpenStudy (anonymous):

lol i know i meant afterwards

OpenStudy (anonymous):

ok i posted my problem

OpenStudy (anonymous):

dx/dy ∫ √ ( t² + 9) [ 0, x ] = √( x² + 9) Then, ( √( x² + 9) )' = 2x/ 2√( x² + 9) = x/ √( x² + 9)